# Deep Multi-Threshold Spiking-UNet for Image Processing

### v2

@@ Abstract
[1]organization=School of Information Science and Technology, University of Science and Technology of China, city=Hefei, citysep=, postcode=230026, country=China
@@ 1 Introduction
U-Net [1] has emerged as a popular choice for a wide range of image-related applications [2, 3, 4, 5], particularly on tasks involving pixel-wise prediction such as image segmentation and denoising. One of the key characteristics of U-Net is its unique use of skip connections, which perform concatenation operations between the corresponding layers of the encoder and the decoder. This strategy allows it to preserve full resolution and maintain contextual semantic information during the feature learning process [6], and to retain the fine-grained details of the target objects. These advantages help U-Net achieve promising performance particularly when limited training samples are available. However, the pixel-wise applications of U-Net on the rapidly emerging neuromorphic devices remain limited.
@@
In recent years, Spiking Neural Networks (SNNs) [7, 8, 9, 10] have emerged as a rapidly growing field of research, driven by their unique ability to emulate the efficient and intricate working scheme of biological nervous systems. In contrast to traditional artificial neural networks, SNNs utilize spiking neurons which convey information through asynchronous, discrete spikes that closely resemble the behavior of biological neurons. This spiking mechanism endows SNNs with inherent advantages such as low energy consumption, fast inference, and event-driven processing [11, 12, 13, 14]. Owing to these benefits, SNNs have attracted substantial interest for their potential in energy-constrained applications such as the Internet of Things (IoT), small satellites, and robotics. By leveraging these bio-inspired spiking mechanisms, SNNs open promising avenues for efficient and high-performance computation across various domains.
@@
Despite the promising advantages of SNNs, implementing them effectively in tasks requiring real-time or rapid inference presents significant challenges. The first of these challenges is the need to maintain efficient information propagation within the network using spikes. To tackle this, researchers have explored the use of ResNet-based deep network topologies [15] and the incorporation of spiking neurons [16, 17, 18]. These advancements have led to the development of deep SNNs, which have shown promising results. The second challenge pertains to the design of effective training strategies. Two primary approaches emerge: converting ANNs to SNNs and training SNNs directly. While these methods have primarily been applied to classification tasks [19, 20, 21], the exploration of SNNs in deep U-Net architectures for pixel-wise tasks remains an under-explored area.
@@
In this paper, we propose the Spiking-UNet, an efficient integration of SNNs and the U-Net architecture for pixel-wise tasks, specifically image segmentation and denoising. To address the challenge of information propagation using spikes, we introduce multi-threshold spiking neurons that fire spikes at different thresholds, enhancing performance in a short time window. This mechanism promotes accurate spike propagation to subsequent layers, ensuring effective information flow. For effective training Spiking-UNet, we construct our model by converting a pre-trained U-Net model and subsequently fine-tuning it. During the ANN-SNN conversion, we observe that the data distribution from different parts of skip connections has significant variability, leading to inconsistent firing rates in Spiking-UNet. To overcome this, we propose a connection-wise normalization strategy, which equalizes the firing rates across skip connections, thereby ensuring more consistent and effective information transmission. In terms of fine-tuning, the traditional Back Propagation Through Time (BPTT) approach commonly used in training SNNs is computationally demanding. To mitigate this, we revise a training method that utilizes an accumulated spiking flow approach to more efficiently update the weights of the converted Spiking-UNet. To validate the effectiveness of our proposed Spiking-UNet, we conduct image segmentation experiments on the DRIVE [22], EM segmentation [23], and CamSeq01 datasets [24], as well as image denoising on the BSD68, and CBSD68 datasets [25]. Experimental results demonstrate that our Spiking-UNet not only exceeds the existing SNN methods but achieves comparable performance to the corresponding U-Net.
@@ 2.1 Spiking Neuron Model
The majority of spiking neural networks predominantly employ the Integrate-and-Fire (IF) model [26] or the Leaky-Integrate-and-Fire (LIF) model [27]. However, these spiking neuron models have shown limitations in efficient information transmission, as highlighted in recent studies [28, 29]. To improve this, Guo et al. [17] employed real-valued spikes (RS) neurons for training and binary spikes for inference. This technique enhances SNNs’ information representation capabilities and demonstrated superior performance on various datasets by the re-parameterization technique. In addition, Lang et al. [16] proposed a multi-level spiking neuron (ML) for improving the efficiency of spiking neural networks. This method enhances gradient propagation and the expression ability of neurons in these networks. As part of our methodology, we propose the multi-threshold spiking neuron, which fires multiple spikes upon reaching different thresholds. This approach aligns the output of the spiking neuron with the activation of the artificial neural network during the conversion process, thereby streamlining the training after conversion.
@@ 2.2 Deep SNN For Image Processing
There are two primary strategies for constructing efficient deep SNNs: one involves converting ANNs to SNNs, and the other entails direct SNN training. In the conversion process, pre-trained ANNs are transformed into SNN counterparts. Diehl et al. [30] proposed a layer-wise weight normalization method combined with threshold balancing to regulate firing rates on image classification task. Likewise, Rueckauer et al. [31] provided some spiking equivalents of common ANN operations and introduced a strategy for selecting robust scale factors for layer-wise normalization. Yan et al. [32] developed a clamped and quantized training method to minimize loss during the conversion process in image classification tasks. Moreover, a recent study in object detection introduced signed neurons with imbalanced thresholds and successfully converted a Tiny YOLO model to an SNN version [33].
Directly training SNNs can be conducted via supervised or unsupervised methods. The classical unsupervised training method for SNNs is Spike-Timing-Dependent-Plasticity (STDP). Diehl et al. [34] applied STDP to train an image classification network. Liu et al. [35] used the STDP method to train an SNN for object recognition with event representation. Liu et al. [36] integrated the STDP method with adaptive thresholding for video-based disguise face recognition. However, supervised methods confront the issue of non-differentiable spiking neuron transfer functions. To address this, Lee et al. [37] proposed a method to smooth spikes and calculate the gradient of each layer using chain rules. Wu et al. [38] presented a supervised method with spatial-temporal backpropagation (STBP) to train SNN models for image classification, leveraging the surrogate gradient. More recently, Chakraborty et al. [39] used a hybrid method combining STDP and STBP to address object detection. Nevertheless, these training methods based on BPTT consume significant memory due to the need to record each timestep’s activation. Kim et al. [40] explored the use of SNN for semantic segmentation, presenting the Spiking-FCN architecture. They showed that direct training Spiking-FCN with surrogate gradient learning yielded better performance and lower latency compared to traditional ANN-to-SNN conversion methods. Wu et al. [41] proposed an efficient spike integration method, ASF-BP, to calculate gradients in image classification, thus reducing resource consumption. In our work, we modify the ASF-BP method to train our SNN model, which substantially reduces the training time for our Spiking-UNet.
@@ 2.3 Combined SNN and U-Net Architecture
The U-Net architecture has found diverse applications within the SNN field. Zhu et al. [42] proposed a potential-assisted spiking neural network combined with U-Net for event-based video reconstruction. Lee et al. [43] presented a hybrid ANN-SNN U-Net architecture, designed to address the issue of spike vanishing. Hagenaars et al. [44] introduced the self-supervised learning of hybrid ANN-SNN U-Net for event-based optical flow. Rançon et al. [45] proposed the novel SNN readout for modified U-Net for depth estimation. Cuadrado et al. [46] combined the 3D encoder U-Net and SNN to extract the temporal context information for optical flow. Patel et al. [47] converted the shallow U-Net to the corresponding spiking version for image segmentation and deployed on a Loihi chip. However, this work only modified the ReLU activation during training U-Net and utilized LIF spiking neurons in the SNN. Compared with Patel’s work, our deep Spiking-UNet utilizes multi-threshold spiking neurons and efficient training after conversion. Besides, our Spiking-UNet solves a potential problem of layer-wise normalization and evaluates across various datasets.
@@ 3.1 Overview
Following the ANN-SNN-Conversion principle, Spiking-UNet adopts the same architecture as the pre-trained U-Net model described in [1]. The configuration of the selected U-Net model closely aligns with the classic setting. The encoder component incorporates four down-sampling stages, while the decoder component utilizes four up-sampling stages. Notably, our baseline model differs from the classic U-Net by employing average pooling operations instead of max pooling operations [31] because average pooling is the form of convolution, which can simulate the relationship between the input and output of the data compared with max pooling. The architecture of the converted Spiking-UNet is depicted in Figure 1(d).
To construct our Spiking-UNet, we first train the U-Net model using conventional back-propagation with supervision labels. Once we have the pre-trained model, we replace the ReLU activation in the ANN with the multi-threshold spiking neuron model. We adjust the weights of kernels and biases in different convolutional layers using weight normalization methods, which will be described in the following subsections. Subsequently, we fine-tune the Spiking-UNet further to reduce the time window. For image segmentation tasks, the segmentation output is obtained by accumulating the spikes through the argmax function in the output layer. In the case of image denoising, the network output is the averaged membrane potential, treated as a noise map. In the first layer of the Spiking-UNet, we adopt static coding with the analog current as the input, following the approach presented in [31, 48]. The analog current remains constant throughout the entire time window. In the subsequent layers, the inputs are spikes fired by neurons in the preceding layer at the previous time step.
@@ 3.3.1 Layer-wise normalization
The layer-wise normalization technique, initially proposed in [30] and extended in [31], employs a scale factor to compute the normalized weights. This scale factor is determined by the ratio between the maximum activation of the previous layer and the maximum activation of the current layer. Mathematically, the normalized parameters are defined as follows:
@@
where $\tilde{W}^{l}$, $W^{l}$, $\lambda^{l}$, and $b^{l}$ denote the normalized weights, pre-trained weights, the maximum activations calculated from the training dataset in ANNs, and bias in the $l^{th}$ layer, respectively. To eliminate outliers, Rueckauer et al. utilized the $99.9\%$ percentile of the maximum activation $\lambda^{l}$ [31], which was also adopted in Spiking-UNet.
@@ 3.4.1 Fine-tuning Principle
The BPTT algorithm is a commonly used training algorithm for SNNs. In addition, recent methods such as STBP [38] and ASF-BP [41] have been proposed. The STBP utilizes the BPTT method through the definition of an approximated derivative, which requires a substantial amount of memory to record the gradient. In our deep Spiking-UNet, the STBP method would cause the GPU memory to overflow. ASF-BP, originally designed for image classification, focuses on training SNNs using the accumulated input and output of an LIF spiking neuron. Compared to STBP, ASF-BP is more computationally efficient as it calculates the gradient only once and requires fewer computational resources.
Motivated by the concept of accumulated spiking flow in ASF-BP [41], we propose a training method for the converted Spiking-UNet model in image segmentation and image denoising tasks. The spiking flow of the multi-threshold spiking neuron is defined as the sum of spikes weighted by the corresponding thresholds. Mathematically, we define the input and output spiking flow as follows:
@@ 4.1.1 Segmentation datasets
To evaluate the performance of Spiking-UNet in image segmentation, we choose three datasets: DRIVE [22], EM segmentation [23], and CamSeq01 [24].
@@ 4.1.2 Denoising datasets
To assess the effectiveness of Spiking-UNet in image denoising, we train the model on the BSD200 dataset [25] and evaluate it on the BSD68 and CBSD68 datasets [25].
@@ 4.2.1 Image Segmentation
Our implementation of the Spiking Fully Convolutional Network (Spiking-FCN) is based on the approach outlined in [40]. We train this model for 100 epochs with a learning rate of $31e-$ 3. We retain all other settings from the original research [40].
@@
We directly train a Spiking-UNet model using multi-level (ML) [16], real spike (RS) [17], and multi-threshold (MT) spiking models. We set the epoch number to 100 and the learning rate to $31e-$ 3, using the Adam optimizer to adjust the learning rate. For the fine-tuning process, we initially use layer-wise and connection-wise weight normalization to adjust weight and bias scales, with the normalization factors of the first layer’s input set to one. In our experiments, we set short time windows to 10, 50, and 20 steps on the three datasets, respectively. After converting from the pre-trained U-Net model, we employ our training method to fine-tune the model. For this fine-tuning phase, we use a learning rate of $1.0\times 10^{-6}$ across all three datasets, and set the epoch number to 100 on the DRIVE, EM, and CamSeq01 datasets.
@@ 4.3.2 Image Segmentation
Dataset T Metric U-Net Spiking-FCN [40] ML [16] RS [17] MT (Ours)
@@
Dataset T Metric $\sigma$ U-Net ML [16] RS [17] MT (Ours)

### v4

@@ Abstract
[1]organization=MoE Key Laboratory of Brain-inspired Intelligent Perception and Cognition, University of Science and Technology of China, city=Hefei, citysep=, postcode=230026, country=China
[…]
[2]organization=Institute of Artificial Intelligence, Hefei Comprehensive National Science Center, city=Hefei, citysep=, postcode=230088, country=China
@@ 1 Introduction
U-Net ronneberger2015u has emerged as a popular choice for a wide range of image-related applications xiao2020global ; zhou2020hierarchical ; liu2022video ; nazir2021ecsu , particularly on tasks involving pixel-wise prediction such as image segmentation and denoising. One of the key characteristics of U-Net is its unique use of skip connections, which perform concatenation operations between the corresponding layers of the encoder and the decoder. This strategy allows it to preserve full resolution and maintain contextual semantic information during the feature learning process cciccek20163d , and to retain the fine-grained details of the target objects. These advantages help U-Net achieve promising performance particularly when limited training samples are available. However, the pixel-wise applications of U-Net on the rapidly emerging neuromorphic devices remain limited.
@@
In recent years, Spiking Neural Networks (SNNs) sengupta2019going ; liao2023convolutional ; zhang2019tdsnn ; yang2022training have emerged as a rapidly growing field of research, driven by their unique ability to emulate the efficient and intricate working scheme of biological nervous systems. In contrast to traditional artificial neural networks, SNNs utilize spiking neurons which convey information through asynchronous, discrete spikes that closely resemble the behavior of biological neurons. This spiking mechanism endows SNNs with inherent advantages such as low energy consumption, fast inference, and event-driven processing akopyan2015truenorth ; davies2018loihi ; pei2019towards ; furber2014spinnaker . Owing to these benefits, SNNs have attracted substantial interest for their potential in energy-constrained applications such as the Internet of Things (IoT), small satellites, and robotics. By leveraging these bio-inspired spiking mechanisms, SNNs open promising avenues for efficient and high-performance computation across various domains.
@@
Despite the promising advantages of SNNs, implementing them effectively in tasks requiring real-time or rapid inference presents significant challenges. The first of these challenges is the need to maintain efficient information propagation within the network using spikes. To tackle this, researchers have explored the use of ResNet-based deep network topologies fang2021deep and the incorporation of spiking neurons feng2022multi ; guo2022real ; fang2021incorporating . These advancements have led to the development of deep SNNs, which have shown promising results. The second challenge pertains to the design of effective training strategies. Two primary approaches emerge: converting ANNs to SNNs and training SNNs directly. While these methods have primarily been applied to classification tasks wu2023dynamic ; xu2023ultra ; chen2022adaptive , the exploration of SNNs in deep U-Net architectures for pixel-wise tasks remains an under-explored area.
@@
In this paper, we propose the Spiking-UNet, an efficient integration of SNNs and the U-Net architecture for pixel-wise tasks, specifically image segmentation and denoising. To address the challenge of information propagation using spikes, we introduce multi-threshold spiking neurons that fire spikes at different thresholds, enhancing performance in a short time window. This mechanism promotes accurate spike propagation to subsequent layers, ensuring effective information flow. For effective training Spiking-UNet, we construct our model by converting a pre-trained U-Net model and subsequently fine-tuning it. During the ANN-SNN conversion, we observe that the data distribution from different parts of skip connections has significant variability, leading to inconsistent firing rates in Spiking-UNet. To overcome this, we propose a connection-wise normalization strategy, which equalizes the firing rates across skip connections, thereby ensuring more consistent and effective information transmission. In terms of fine-tuning, the traditional Back Propagation Through Time (BPTT) approach commonly used in training SNNs is computationally demanding. To mitigate this, we revise a training method that utilizes an accumulated spiking flow approach to more efficiently update the weights of the converted Spiking-UNet. To validate the effectiveness of our proposed Spiking-UNet, we conduct image segmentation experiments on the DRIVE staal2004ridge , EM segmentation cardona2010integrated , and CamSeq01 datasets fauqueur2007assisted , as well as image denoising on the BSD68, and CBSD68 datasets martin2001database . Experimental results demonstrate that our Spiking-UNet not only exceeds the existing SNN methods but achieves comparable performance to the corresponding U-Net.
@@ 2.1 Spiking Neuron Model
The majority of spiking neural networks predominantly employ the Integrate-and-Fire (IF) model abbott1999lapicque or the Leaky-Integrate-and-Fire (LIF) model doutsi2021dynamic . However, these spiking neuron models have shown limitations in efficient information transmission, as highlighted in recent studies li2022brain ; wu2021liaf . To improve this, Guo et al. guo2022real employed real-valued spikes (RS) neurons for training and binary spikes for inference. This technique enhances SNNs’ information representation capabilities and demonstrated superior performance on various datasets by the re-parameterization technique. In addition, Lang et al. feng2022multi proposed a multi-level spiking neuron (ML) for improving the efficiency of spiking neural networks. This method enhances gradient propagation and the expression ability of neurons in these networks. As part of our methodology, we propose the multi-threshold spiking neuron, which fires multiple spikes upon reaching different thresholds. This approach aligns the output of the spiking neuron with the activation of the artificial neural network during the conversion process, thereby streamlining the training after conversion.
@@ 2.2 Deep SNN For Image Processing
There are two primary strategies for constructing efficient deep SNNs: one involves converting ANNs to SNNs, and the other entails direct SNN training. In the conversion process, pre-trained ANNs are transformed into SNN counterparts. Diehl et al. diehl2015fast proposed a layer-wise weight normalization method combined with threshold balancing to regulate firing rates on image classification task. Likewise, Rueckauer et al. rueckauer2017conversion provided some spiking equivalents of common ANN operations and introduced a strategy for selecting robust scale factors for layer-wise normalization. Yan et al. yan2021near developed a clamped and quantized training method to minimize loss during the conversion process in image classification tasks. Moreover, a recent study in object detection introduced signed neurons with imbalanced thresholds and successfully converted a Tiny YOLO model to an SNN version kim2018spiking .
Directly training SNNs can be conducted via supervised or unsupervised methods. The classical unsupervised training method for SNNs is Spike-Timing-Dependent-Plasticity (STDP). Diehl et al. diehl2015unsupervised applied STDP to train an image classification network. Liu et al. liu2020unsupervised used the STDP method to train an SNN for object recognition with event representation. Liu et al. liu2019deep integrated the STDP method with adaptive thresholding for video-based disguise face recognition. However, supervised methods confront the issue of non-differentiable spiking neuron transfer functions. To address this, Lee et al. lee2016training proposed a method to smooth spikes and calculate the gradient of each layer using chain rules. Wu et al. wu2018spatio presented a supervised method with spatial-temporal backpropagation (STBP) to train SNN models for image classification, leveraging the surrogate gradient. More recently, Chakraborty et al. chakraborty2021fully used a hybrid method combining STDP and STBP to address object detection. Nevertheless, these training methods based on BPTT consume significant memory due to the need to record each timestep’s activation. Kim et al. kim2022beyond explored the use of SNN for semantic segmentation, presenting the Spiking-FCN architecture. They showed that direct training Spiking-FCN with surrogate gradient learning yielded better performance and lower latency compared to traditional ANN-to-SNN conversion methods. Wu et al. wu2021training proposed an efficient spike integration method, ASF-BP, to calculate gradients in image classification, thus reducing resource consumption. In our work, we modify the ASF-BP method to train our SNN model, which substantially reduces the training time for our Spiking-UNet.
@@ 2.3 Combined SNN and U-Net Architecture
The U-Net architecture has found diverse applications within the SNN field. Zhu et al. zhu2022event proposed a potential-assisted spiking neural network combined with U-Net for event-based video reconstruction. Lee et al. lee2020spike presented a hybrid ANN-SNN U-Net architecture, designed to address the issue of spike vanishing. Hagenaars et al. hagenaars2021self introduced the self-supervised learning of hybrid ANN-SNN U-Net for event-based optical flow. Rançon et al. ranccon2022stereospike proposed the novel SNN readout for modified U-Net for depth estimation. Cuadrado et al. cuadrado2023optical combined the 3D encoder U-Net and SNN to extract the temporal context information for optical flow. Patel et al. patel2021spiking converted the shallow U-Net to the corresponding spiking version for image segmentation and deployed on a Loihi chip. However, this work only modified the ReLU activation during training U-Net and utilized LIF spiking neurons in the SNN. Compared with Patel’s work, our deep Spiking-UNet utilizes multi-threshold spiking neurons and efficient training after conversion. Besides, our Spiking-UNet solves a potential problem of layer-wise normalization and evaluates across various datasets.
@@ 3.1 Overview
Following the ANN-SNN-Conversion principle, Spiking-UNet adopts the same architecture as the pre-trained U-Net model described in ronneberger2015u . The configuration of the selected U-Net model closely aligns with the classic setting. The encoder component incorporates four down-sampling stages, while the decoder component utilizes four up-sampling stages. Notably, our baseline model differs from the classic U-Net by employing average pooling operations instead of max pooling operations rueckauer2017conversion because average pooling is the form of convolution, which can simulate the relationship between the input and output of the data compared with max pooling. The architecture of the converted Spiking-UNet is depicted in Figure 1(d).
To construct our Spiking-UNet, we first train the U-Net model using conventional back-propagation with supervision labels. Once we have the pre-trained model, we replace the ReLU activation in the ANN with the multi-threshold spiking neuron model. We adjust the weights of kernels and biases in different convolutional layers using weight normalization methods, which will be described in the following subsections. Subsequently, we fine-tune the Spiking-UNet further to reduce the time window. For image segmentation tasks, the segmentation output is obtained by accumulating the spikes through the argmax function in the output layer. In the case of image denoising, the network output is the averaged membrane potential, treated as a noise map. In the first layer of the Spiking-UNet, we adopt static coding with the analog current as the input, following the approach presented in rueckauer2017conversion ; kim2020spiking . The analog current remains constant throughout the entire time window. In the subsequent layers, the inputs are spikes fired by neurons in the preceding layer at the previous time step.
@@ 3.3.1 Layer-wise normalization
The layer-wise normalization technique, initially proposed in diehl2015fast and extended in rueckauer2017conversion , employs a scale factor to compute the normalized weights. This scale factor is determined by the ratio between the maximum activation of the previous layer and the maximum activation of the current layer. Mathematically, the normalized parameters are defined as follows:
@@
where $\tilde{W}^{l}$, $W^{l}$, $\lambda^{l}$, and $b^{l}$ denote the normalized weights, pre-trained weights, the maximum activations calculated from the training dataset in ANNs, and bias in the $l^{th}$ layer, respectively. To eliminate outliers, Rueckauer et al. utilized the $99.9\%$ percentile of the maximum activation $\lambda^{l}$ rueckauer2017conversion , which was also adopted in Spiking-UNet.
@@ 3.4.1 Fine-tuning Principle
The BPTT algorithm is a commonly used training algorithm for SNNs. In addition, recent methods such as STBP wu2018spatio and ASF-BP wu2021training have been proposed. The STBP utilizes the BPTT method through the definition of an approximated derivative, which requires a substantial amount of memory to record the gradient. In our deep Spiking-UNet, the STBP method would cause the GPU memory to overflow. ASF-BP, originally designed for image classification, focuses on training SNNs using the accumulated input and output of an LIF spiking neuron. Compared to STBP, ASF-BP is more computationally efficient as it calculates the gradient only once and requires fewer computational resources.
Motivated by the concept of accumulated spiking flow in ASF-BP wu2021training , we propose a training method for the converted Spiking-UNet model in image segmentation and image denoising tasks. The spiking flow of the multi-threshold spiking neuron is defined as the sum of spikes weighted by the corresponding thresholds. Mathematically, we define the input and output spiking flow as follows:
@@ 4.1.1 Segmentation datasets
To evaluate the performance of Spiking-UNet in image segmentation, we choose three datasets: DRIVE staal2004ridge , EM segmentation cardona2010integrated , and CamSeq01 fauqueur2007assisted .
@@ 4.1.2 Denoising datasets
To assess the effectiveness of Spiking-UNet in image denoising, we train the model on the BSD200 dataset martin2001database and evaluate it on the BSD68 and CBSD68 datasets martin2001database .
@@ 4.2.1 Image Segmentation
Our implementation of the Spiking Fully Convolutional Network (Spiking-FCN) is based on the approach outlined in kim2022beyond . We train this model for 100 epochs with a learning rate of $31e-$ 3. We retain all other settings from the original research kim2022beyond .
@@
We directly train a Spiking-UNet model using multi-level (ML) feng2022multi , real spike (RS) guo2022real , and multi-threshold (MT) spiking models. We set the epoch number to 100 and the learning rate to $31e-$ 3, using the Adam optimizer to adjust the learning rate. For the fine-tuning process, we initially use layer-wise and connection-wise weight normalization to adjust weight and bias scales, with the normalization factors of the first layer’s input set to one. In our experiments, we set short time windows to 10, 50, and 20 steps on the three datasets, respectively. After converting from the pre-trained U-Net model, we employ our training method to fine-tune the model. For this fine-tuning phase, we use a learning rate of $1.0\times 10^{-6}$ across all three datasets, and set the epoch number to 100 on the DRIVE, EM, and CamSeq01 datasets.
@@ 4.3.2 Image Segmentation
Dataset T Metric U-Net Spiking-FCN kim2022beyond ML feng2022multi RS guo2022real MT (Ours)
@@
Dataset T Metric $\sigma$ U-Net ML feng2022multi RS guo2022real MT (Ours)